OpenStudy (anonymous):

Anyone please help w/ steps to simplify ᵌ√729x^4y^5 (everythng is under the radical)

OpenStudy (dumbcow):

break it into parts \[\sqrt[3]{729}=9\] \[\sqrt[3]{x^{4}} = \sqrt[3]{x ^{3}x ^{1}} = x \sqrt[3]{x}\] \[\sqrt[3]{y^{5}}=\sqrt[3]{y^{3}y^{2}}=y \sqrt[3]{y^{2}}\]

OpenStudy (anonymous):

9xy ᵌ√xy^2

OpenStudy (anonymous):

thank you so much